### v2

@@ 5.1 Definition of a cactus
  • At least one element of $s$ has depth exactly $n-1$.
@@
[…]
Figure 8: A graph $G$ and a cactus of depth $3$ on $G$. Every subtree is a cactus $C$ whose root is labeled by $\xi(C)$, whose children are $s(C)$ (ordered from bottom to top) and where the edge to each child is labeled by the corresponding $\chi_{i}(C)$.
@@ 5.3 Encoding cycles by cactus forests of bounded depth
The cycles $d^{(j)}$ can be seen as the maximal cycles which appear in $c$ in which $c_{0}$ does not appear. Since the cycles $d^{(1)},\ldots,d^{(k)}$ are on the subgraph $H^{\prime}$ of $H$ on vertices $V_{H}\backslash\{c_{0}\}$, by induction each of these cycles is encoded by a cactus forest of depth no larger than $|V_{H}|-1$. This implies that $c$ is encoded by a cactus forest of depth no larger than $|V_{H}|$.∎
@@ 6.1 Transforming cycles with additional spines
There exists a constant $\lambda_{G}$ such that, for any cycle $c$ which is decomposable into squares and spine $t$ on $c_{0}$, we have:
[…]
@@
Let $c$ be a cycle which is decomposable into squares and let $t$ be a spine on $c_{0}$. We have:
@@ 6.2 Parallelization
This derives from the fact that for all $i$, $r^{(i)}\mathcal{R}_{\epsilon}r^{(i)}$ and that the second letter and the penultimate letter of $r^{(i)}$ are both neighbors of $a$. ∎
@@
[…]
@@
For any $j$, as a consequence of the hypothesis, the difference $\sum_{i\leq j}\textbf{z}^{\prime}_{i}-\sum_{i\leq j}\textbf{z}_{i}$ is equal to 00 or $2$. Whenever $\sum_{i\leq j}\textbf{z}^{\prime}_{i}-\sum_{i\leq j}\textbf{z}_{i}=2$, since
@@
Let us fix $\textbf{r}=(r^{(j)})_{0\leq j\leq l}$ a sequence of cycles all beginning and ending at some vertex $a\in V_{G}$. Let $c$ be a cycle decomposable into squares and $t$ a spine on $c_{0}$. For all $n\geq 0$, we have
@@
    Let us prove that there are sequences $\textbf{z}^{k}$, $k\leq n$, such that $^{0}=\textbf{z}$ z and
@@
    From $\gamma_{k-1}$ to $\gamma_{k}$ moving spines: Let us assume that $\textbf{z}^{0},\dots,\textbf{z}^{k-1}$ have been defined, where $k\geq 1$. Let the sequence of cycles $(\gamma_{k,j})_{j\in\llbracket 0,2\lambda_{G}\rrbracket}$ be defined for all $k$ and $j$ as the concatenation of $2^{k-1}$ identical words that we call blocks:
@@
    In the definition of $\Gamma^{\textbf{r}}_{\textbf{z}^{k,j}}(\gamma_{k,j})$ it may happen that $\Gamma^{\textbf{r}}$ inserts an element of r in between $t^{\lambda_{G}l(c)/2}$ and $t^{j}$, which changes the relative position of the spine being moved in the $i$-th block when constructing $\gamma_{k,j+1}$ from $\gamma_{k,j}$. To cancel out this movement, we define a quantity $\omega^{k,j+1}_{q}$ to be $1$ when this happens, and 00 otherwise, and we define:
@@
    $\underbrace{\left(\lambda_{G}l(c)-2j\right)+\overbrace{i\cdot\lambda_{G}2^{n-k+1}l(c)}^{\text{the spine is in block $i+1$}}}_{\text{position of spine $t$ before moving}}<\underbrace{\textbf{z}^{k,j}_{0}+\ldots+\textbf{z}^{k,j}_{q}}_{\text{insertion position of $r^{(q)}$}}$
@@
      • $\textbf{x}^{p+1}_{q}=\textbf{x}^{p}_{q}-2$ if $q$ is such an index but $q-1$ is not (or $q=0$);
      • $\textbf{x}^{p+1}_{q}=\textbf{x}^{p}_{q}+2$ if $q$ is not such an index but $q-1$ is;
@@ 6.3 The gap function is logarithmic
Let $t^{(i)}$ be the spine $u_{i}u_{i+1}u_{i}$ and $u^{(i)}=u_{0}\dots u_{i}\odot\left(t^{(i)}\right)^{m}\odot u_{i}\dots u_{l(u)}$. It is enough to check that $(t^{m}\odot u)\mathcal{R}_{0}u^{(1)}$, that $u^{(i)}\mathcal{R}_{0}u^{(i+1)}$ for all $10<i<l(u)-$ 1, and that $u^{(l(u)-1)}\mathcal{R}_{0}(u\odot t^{\prime m})$. ∎
@@
[…]
@@
    Gathering the spines: Now we bring every spine at depth $k$ to depth $k-1$. We consider two cases:
@@
    By repeating this argument, and since $n(C_{1},\ldots,C_{l})\leq|V_{G}|$, we find a cycle $f_{q}$ which is represented by a cactus forest of depth $1$ such that

### v3

@@ 5.1 Definition of a cactus
  • At least one element of $s$ has depth exactly n1n-1.
@@
abc
1012abcaabcaaaaabcabbcabcaccacabcaabca
Figure 8: A graph $G$ and a cactus of depth 33 on $G$. Every subtree is a cactus $C$ whose root is labeled by $\xi(C)$, whose children are $s(C)$ (ordered from bottom to top) and where the edge to each child is labeled by the corresponding $\chi_{i}(C)$.
@@ 5.3 Encoding cycles by cactus forests of bounded depth
[…]
@@ 6.1 Transforming cycles with additional spines
There exists a constant $\lambda_{G}$ such that, for any simple cycle $c$ which is decomposable into squares and spine $t$ on $c_{0}$, we have:
$d^{\mathcal{R}}_{G}(t^{\lambda_{G}}\odot c,t^{\lambda_{G}+l(c)/2})\leq\lambda_{G}\qquad\text{and}\qquad l(c)\leq 2\lambda_{G}.$
@@
Let $c$ be a simple cycle which is decomposable into squares and let $t$ be a spine on $c_{0}$. We have:
@@ 6.2 Parallelization
This derives from the fact that $r^{(i)}\mathcal{R}_{\epsilon}r^{(i)}$ for all $i$ and $\epsilon$, and that the second letter and the penultimate letter of $r^{(i)}$ are both neighbors of $a$. ∎
@@
Let us fix $\textbf{r}=(r^{(j)})_{0\leq j\leq l}$ a sequence of cycles all beginning and ending at some vertex $a\in V_{G}$. Let us also consider a pair $(\textbf{z}$, $\textbf{z}^{\prime})$ for which there exists an increasing sequence $(j_{i})_{1\leq i\leq 2\tau}$ of integers in $\llbracket 0,l\rrbracket$ such that:
@@
For any $j$, as a consequence of the hypothesis, the difference $\sum_{i\leq j}\textbf{z}^{\prime}_{i}-\sum_{i\leq j}\textbf{z}_{i}$ is equal to 0 or $2$. Whenever $\sum_{i\leq j}\textbf{z}^{\prime}_{i}-\sum_{i\leq j}\textbf{z}_{i}=2$, since
@@
Let us fix $\textbf{r}=(r^{(j)})_{0\leq j\leq l}$ a sequence of cycles all beginning and ending at some vertex $a\in V_{G}$. Let $c$ be a simple cycle decomposable into squares and $t$ a spine on $c_{0}$. For all $n\geq 0$, we have
@@
    Let us prove that there are sequences $\textbf{z}^{k}$, $k\leq n$, such that $\textbf{z}^{0}=\textbf{z}$ and
@@
    From $\gamma_{k-1}$ to $\gamma_{k}$ - moving spines: Let us assume that $\textbf{z}^{0},\dots,\textbf{z}^{k-1}$ have been defined, where $k\geq 1$. Let the sequence of cycles $(\gamma_{k,j})_{j\in\llbracket 0,2\lambda_{G}\rrbracket}$ be defined for all $k$ and $j$ as the concatenation of $2^{k-1}$ identical words that we call blocks:
@@
    In the definition of $\Gamma^{\textbf{r}}_{\textbf{z}^{k,j}}(\gamma_{k,j})$ it may happen that $\Gamma^{\textbf{r}}$ inserts an element of r in between $t^{\lambda_{G}l(c)/2}$ and $t^{j}$, which changes the relative position of the spine being moved in the $i$-th block when constructing $\gamma_{k,j+1}$ from $\gamma_{k,j}$. To cancel out this movement, we define a quantity $\omega^{k,j+1}_{q}$ to be $1$ when this happens, and 0 otherwise, and we define:
@@
[…]
@@
      • $\textbf{x}^{p+1}_{q}=\textbf{x}^{p}_{q}-2$ if $q$ is such an index but q1q-1 is not (or $q=0$);
      • $\textbf{x}^{p+1}_{q}=\textbf{x}^{p}_{q}+2$ if $q$ is not such an index but q1q-1 is;
@@ 6.3 The gap function is logarithmic
Let $t^{(i)}$ be the spine $u_{i}u_{i+1}u_{i}$ and $u^{(i)}=u_{0}\dots u_{i}\odot\left(t^{(i)}\right)^{m}\odot u_{i}\dots u_{l(u)}$. It is enough to check that $(t^{m}\odot u)\mathcal{R}_{0}u^{(1)}$, that $u^{(i)}\mathcal{R}_{0}u^{(i+1)}$ for all 0<i<l(u)10<i<l(u)-1, and that $u^{(l(u)-1)}\mathcal{R}_{0}(u\odot t^{\prime m})$. ∎
@@
$d_{G}(r\odot c\odot r^{\prime},r\odot c^{\prime}\odot r^{\prime})\leq d_{G}(c,c^{\prime})+4\text{diam}(G).$
@@
    Gathering the spines: Now we bring every spine at depth $k$ to depth k1k-1. We consider two cases:
@@
    By repeating this argument, and since $n(C_{1},\ldots,C_{l})\leq|V_{G}|$, we find a cycle $f$ which is represented by a cactus forest of depth $1$ such that